OpenStudy (anonymous):

Please help someone:) describe how to solve a linear equation using the equality properties. Demonstrate the process with an example. Next, replace the equal sign in your example with an inequality by using the less than or greater than sign. Then solve the inequality.

OpenStudy (anonymous):

If A = B ... and ... A = C ... then ... B = C By the equality property if one side of two or more equations equal the same thing, then the other sides equal each other. If y = x - 1 y = 3x - 7 then x - 1 = 3x - 7 2x = 6 x = 3 note this can be extended to non-linear equations as well if y^3 = x - 1 and y^3 = x - 7 then x still has to equal 3